OpenStudy (anonymous):

Write the quadratic function in the form f(x)=a−(x-h)^ 2+k . Then, give the vertex of its graph.f(x)=-2x^2+8x-5

OpenStudy (zzr0ck3r):

\(f(x) = ax^2+bx+c\) is the same thing as \(f(x) =a(x+\dfrac{b}{2a})^2+c-a(\dfrac{b}{2a})^2 \)

OpenStudy (anonymous):

so i just plug it in?

OpenStudy (sepeario):

isn't the x point for the vertex -b/2a?

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

this is not that hard it is \[f(x)=-2x^2+8x-5\]right?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

you want to write it as \[f(x)=a(x-h)^2+k\] all you need is \(h\) and \(k\)

OpenStudy (misty1212):

since evidently \(a=-2\)

OpenStudy (misty1212):

to find \(h\) compute \(-\frac{b}{2a}\) which, in your case, is \(-\frac{8}{2\times (-2)}\)

OpenStudy (misty1212):

what do you get ?

OpenStudy (anonymous):

is it -8/4x? and simplify so its -2x?

OpenStudy (anonymous):

or is it 2/x

OpenStudy (anonymous):

it's just 2 Kimes

OpenStudy (anonymous):

oh opps i mistaken x and a variable

OpenStudy (anonymous):

as*

OpenStudy (zzr0ck3r):

@Kimes yes just plug things :)